OpenStudy (anonymous):

5(2n+1)+ -5= 10 true or false

OpenStudy (igreen):

Solve: \(5(2n + 1) + -5 = 10\) \(5(2n + 1) - 5 = 10\) \(10n + 5 - 5 = 10\) \(10n = 10\) \(n = 1\)

OpenStudy (anonymous):

Not quite sure what you mean by true or false but if you solve it you get an answer. Distribute 10n+5+(-5)=10 Positive 5 and negative 5 cancel out. 10n=10 n=1

OpenStudy (igreen):

What's with the "True or False"? @wakaflacka

OpenStudy (anonymous):

for what value of n?

OpenStudy (anonymous):

@iGreen He had answered the question; he was asking if he got it right or not.

OpenStudy (igreen):

Oh, I didn't know.

OpenStudy (anonymous):

so its false

OpenStudy (anonymous):

thx :)